OpenStudy (anonymous):

what is the equivalent weight of h2so4 in the reaction H2SO4+NAI = NA2SO4+I2+H2S+H2O

OpenStudy (anonymous):

H2SO4 molecular weight = 2+32+64= 98 equivalent weight= molecular wt / no. of e donation = 98/2= 49